#1)
-7x--26+6x
#2
-x+12=-4+x

Answer by xcentaur(357)

#1)
-7x--26+6x
You've put two '-' signs, and theres no equal to. This expression cannot be simplified as I can't figure out what to do with the double minus signs, nor can I solve it because there is no equal to sign.Please check your question and post again.


#2
-x+12=-4+x
-x-x=-4-12
-2x=-16
x=8


Hope this helps,
prabhat